are the following perpendicular, parallel or neither? How do you know? a) 9x-12y=17 and 3x-4y= 17 b) 5x+6y=30 and 6x+5y=24
Answer #1

We know that two lines angle between them is oo & if angle between them is are are 90o. parallel if perpendiculer If y=mx+c &and if lines are perpendicular m.-m2 ic 0-90° then tango = tm, mz o = m m₂ m, m2 It m, mz 1+m, m2=0 mm, =-17 (a) Given lines